OpenStudy (anonymous):

@jessiehuang_ Evaluate 5(a – 2b) when a = –3 and b = 5. A. 65 B. 35 C. –65 D. –25

OpenStudy (anonymous):

@jessiehuang_

OpenStudy (anonymous):

@jessiehuang_

OpenStudy (anonymous):

-65

OpenStudy (anonymous):

5(a - 2b) a = -3, b = 5 substitute your values of a and b into your expression 5(-3 - 2(5)) 5(-3-10) 5(-13) -65

OpenStudy (anonymous):

Evaluate –6(x – 2y) when x = –3 and y = –5 A. –42 B. 80 C. –80 D. –16

OpenStudy (anonymous):

-42

OpenStudy (anonymous):

these are not that hard to do...just sub in your values into the expression and solve mathematically. jessie..be careful just giving answer...you can get booted off for that

OpenStudy (anonymous):

Simplify: (–5) • (7) – (6) • (–2) A. –23 B. –35 C. 10 D. 47

OpenStudy (anonymous):

oh i'm sorry, i'm multi-tasking with my english homework and open study, i'll try and show explanations from now on @kelliegirl33 but thanks for the heads up :)

OpenStudy (anonymous):

Simplify: (–5) • (7) – (6) • (–2) A. –23 B. –35 C. 10 D. 47

OpenStudy (anonymous):

no problem...I learned the hard way and got booted for 2 weeks by just giving answers

OpenStudy (anonymous):

Simplify: (–5) • (7) – (6) • (–2) A. –23 B. –35 C. 10 D. 47

OpenStudy (anonymous):

ok just multiply from left to right; -5*7=-35 -(-6)*(-2)=12 combine: -23

OpenStudy (anonymous):

so its A

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Simplify: –5(x – 3) + 2(a – 9) A. –5x + 2a + 33 B. –5x + 2a + 4 C. –5x + 2a – 3 D. –5x + 2a – 33

OpenStudy (anonymous):

ok so distribute both parenthesis: -5x+15+2a-18 rearrage: -5x+2a+15-18 combine like terms: -5x+2a-3 so c

OpenStudy (anonymous):

Simplify: –6(x – 3) – 4(y + 5) + (–5)(7 – 2b) A. –6x –4y – 10b – 3 B. –6x –4y + 10b – 37 C. –6x –4y + 10b + 3 D. –6x –4y + 10b + 33

OpenStudy (anonymous):

–6(x – 3) – 4(y + 5) + (–5)(7 – 2b) distribute: -6x+18-4y-20-35+10b rearrange: -6x-4y+10b+18-20-35 combine like terms: -6x-4y+10b-37 b

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

u got me 100% again omg u should be a mod

OpenStudy (anonymous):

close and reopen ok

OpenStudy (anonymous):

just make sure to tag.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Which equation could be used to solve this problem? The sum of two consecutive integers is 141. Find the integers. A. n + 1= 141 B. 2n + 2 = 141 C. 2n = 141 D. 2n + 1 = 141